# Beyond Pixels: Enhancing LIME with Hierarchical Features and Segmentation Foundation Models

### v4

@@ 5.3 Qualitative evaluation
User study. Following the methodology by Chromik & Schuessler (2020), we conducted a user study (approved by the institute’s ethics council) to assess the interpretability of the explanations. This study involved 87 participants recruited via Amazon Mechanical Turk (MTurk) and included 20 randomly of the 100 images in our dataset (Section D.1). These images were accompanied by explanations using DSEG and other segmentation techniques within the LIME framework (Section 5.1).
@@
Participants rated the explanations on a scale from 1 (least effective) to 5 (most effective) based on their intuitive understanding and the predicted class. Table 3 summarizes the average scores, the cumulative number of top-rated explanations per instance, and the statistical significance of user study results for each segmentation approach. DSEG is most frequently rated as the best and consistently ranks high even when it is not the leading explanation. Paired t-tests indicate that DSEG is statistically significantly superior (additional results in Section D.2).
@@ 5.4 Limitations and future work
DSEG-LIME performs the feature generation directly on images before inputting them into the model for explanation. For models like ResNet with smaller input sizes (He et al., 2015), the quantitative advantages of DSEG are less evident (Section B.1). Experiments have shown that better results can be achieved with a lower stability score threshold of SAM. Furthermore, substituting superpixels with a specific value in preservation and deletion evaluations can introduce an inductive bias (Garreau & Mardaoui, 2021). To reduce this bias, using a generative model to synthesize replacement areas could offer a more neutral alteration. Additionally, future work should thoroughly evaluate feature attribution maps to ensure that methods assign significant attributions to the correct regions, like in Section C.6. This comprehensive assessment is essential for verifying the interpretability and reliability of such methods. Lastly, our approach, like any other LIME-based method (Ribeiro et al., 2016; Zhou et al., 2021; Zhao et al., 2020; Tan et al., 2024), does not assume a perfect match between the explanation domains and the model’s actual domains since it simplifies the model by a local surrogate. Nonetheless, our quantitative analysis confirms that the approximations closely reflect the model’s behaviour. Future work could focus on integrating the foundation model directly into the system through a model-intrinsic approach, similar to (Sun et al., 2023).
No free lunch. Although DSEG provides promising results in many domains, it is not always universally applicable. When domain-specific knowledge is crucial to identify meaningful features or the feature generation task is inherently complex, DSEG might not perform as effectively as traditional segmentation methods within LIME (Khani et al., 2024) (Section C.5). However, future exploration could involve testing alternative segmentation techniques, such as integrating HSSN (Wang et al., 2023) or HIPPIE (Li et al., 2022) instead of SAM (or DETR) to overcome this limitation.
@@ 6 Conclusion
In this study, we introduced DSEG-LIME, an extension to the LIME framework, incorporating segmentation foundation models for feature generation with hierarchical feature calculation. This approach ensures that the generated features more accurately reflect human-recognizable concepts, enhancing the interpretability of explanations. Furthermore, we refined the process of feature attribution within LIME through an iterative method, establishing a segmentation hierarchy that contains the relationships between components and their subcomponents. In Section C.2, we show that our idea also helps explain a model’s wrong classifications. Through a comprehensive two-part evaluation, split into quantitative and qualitative analysis, DSEG emerged as the superior method, outperforming other LIME-based approaches in most evaluated metrics. The adoption of foundational models marks a significant step towards enhancing the post-hoc and model-agnostic interpretability of deep learning models.
@@ A.1 DSEG-LIME algorithm
1:  Input: $f$ (black-box model), $\zeta$ (segmentation function), $x$ (input instance), $g$ (interpretable model), $d$ (maximum depth), $hp$ (segmentation hyperparameters), $\theta$ (minimum segment size), $k$ (top segments to select)
@@
5:  $\mathcal{S^{\prime}}\leftarrow\{s_{i}\in\mathcal{S}\,|\,\text{size}(s_{i})\geq\theta\}$
@@
7:  $\mathcal{H}\leftarrow\text{BuildHierarchy}(\mathcal{S^{\prime}})$
@@
First, we identify the top-level segments, which do not occur as subparts of any other segments. These segments serve as the highest-level nodes in the hierarchical graph. Starting from these top-level segments, we apply a top-down approach to identify child segments recursively. For each parent segment, we check for segments within it; these segments are designated as child nodes of the parent in the graph.
@@ A.2 Hierarchical Composition
The algorithm, designed for hierarchical composition, constructs a graph to represent the segments of a given instance, as shown in Algorithm 2. A key feature of this algorithm is the additional hyperparameter $t$, which determines when one segment is considered a subpart of another based on their overlap. Initially, the graph includes loops and redundant edges. To ensure it accurately represents the hierarchical relationships, the algorithm applies a series of predefined rules to tune the graph, removing unnecessary connections and simplifying its structure for the final representation.
@@ B.1 ResNet
In Table 4, we detail the quantitative results for the ResNet-101 model, comparing our evaluation with the criteria used for EfficientNetB4 under consistent hyperparameter settings. The review encompasses a comparative analysis with EfficientNetB3, emphasizing performance under contrastive conditions. The findings substantiate the results obtained from EfficientNet, demonstrating that the LIME techniques exhibit unpredictable behavior in the presence of model noise or prediction shuffling, despite varied segmentation strategies. This suggests an inherent randomness in the model explanations. Furthermore, the results indicate that all XAI methodologies displayed performance levels that were inferior to those of EfficientNetB4. However, it is significant to note that DSEG surpassed all other methods in terms of performance.
@@
Table 5 presents further findings of ResNet. SLIME with DSEG yields the lowest AUC for incremental deletion, whereas Quickshift and Felzenszwalb show the highest. WS produces the smallest superpixels for compactness, contrasting with DSEG’s larger ones. The stability analysis shows that all segmentations are almost at the same level, with QS being the best and GLIME the best-performing overall. Echoing EfficientNet’s review, segmentation defines runtime, with DSEG being the most time-consuming.
@@ B.2 VisionTransformer
Table 6 provides the quantitative results for the VisionTransformer (ViT-384) model, employing settings identical to those used for EfficientNet and ResNet, with ViT processing input sizes of (384x384). The class-specific results within this table align closely with the performances recorded for the other models, further underscoring the effectiveness of DSEG. This consistency in DSEG performance is also evident in the data presented in Table 7.
@@ B.3 ConvNext
Table 8 and Table 9 present the quantitative results pertaining to the ConvNext model (Liu et al., 2022), which is the last model that has been thoroughly evaluated in this study. The settings utilized in this evaluation remain consistent with those employed previously.
@@ B.4 DSEG compared to EAC
We conducted additional experiments with Explain Any Concept (EAC) (Sun et al., 2023), performing the same quantitative experiments as for DSEG, but with a reduced dataset comprising 50 images to optimize computational time (Section D.1). We began our evaluation by noting that EAC, unlike DSEG-LIME, cannot be applied to arbitrary models, which is a significant drawback of their method and prevents comprehensive comparisons. Thus, we compared our approach against LIME and EAC, in explaining ResNet. The results are listed in Table 10 and Table 11.
@@ B.5 DSEG within SLICE
Table 12 reports the results for the DSEG framework applied to the SLICE (Bora et al., 2024) approach and conducts a comparative evaluation against Quickshift (QS) in LIME. The experiments were carried out on the minimized dataset consisting of 50 images in total. Each image was processed for 200 steps and repeated 8 times. The primary evaluation metric used was the average rank similarity (p = 0.3), which focused on positive and negative segments and ensured comparability by employing the minimum segment count of both approaches. Furthermore, a sign flip analysis was performed, measuring the number of segments that changed signs during the computation. This analysis considered all segments produced by each technique.
@@ B.6 DETR within DSEG
In Table 13 and Table 14, we conducted the DETR experiments within LIME. Based on prior results, we assess its performance by contrasting it with SLIC within the LIME framework, also utilizing the same condensed dataset comprising 50 images. Both experiments were configured with identical parameters, and DETR was implemented for basic panoptic segmentation.
Table 13: Quantitative summary - classes DETR. The table showcases metrics for EfficientNetB4, specifically at a finer concept granularity; the hierarchical segmentation tree has a depth of two. Results reported pertain solely to integrating DSEG and SLIC within the scope of the LIME frameworks examined.
@@
Table 14: Quantitative summary - numbers DETR. The table showcases the numeric values in the same manner as in Table 13 but for numeric values.
@@
DETR demonstrates superior performance on the dataset compared to the LIME variants utilizing SLIC. Despite its efficacy, the segmentation quality of DETR was generally inferior to that of SAM, as evidenced by less compact explanations. This observation is further supported by the examples in Figure 4. The visualizations reveal that DETR often segments images in ways that do not align with typical human-recognizable concepts, highlighting a potential limitation in its practical utility for generating explanatory segments. Moreover, DETR does not support the construction of a segmentation hierarchy, lacking the ability to produce finer and coarser segments, which diminishes its flexibility compared to methods such as SAM.
@@
B.7 DSEG-LIME with SAM 2
In the main paper, we conducted experiments using SAM 1. In this part, we integrate SAM 2 (Ravi et al., 2024) with the ’hiera_l’ backbone into the DSEG framework, applying a 0.8 stability score threshold. We also report the results for EfficientNetB4 on the dataset comprising 50 images in Table 15 and Table 16.
@@ B.7 DSEG-LIME with SAM 2
Exemplary explanations. Figure 5 presents explanations generated by DSEG using both SAM 1 and SAM 2, highlighting cases where the newer version of SAM enables DSEG to produce more meaningful and interpretable explanations. Each image includes explanations for the predicted class from EfficientNetB4. While SAM 2 shows improved segmentation in these examples, similar results can be obtained with SAM 1 by appropriately adjusting the hyperparameters for automatic mask generation.
@@ B.8 EfficientNetB4 with depth of two
In Table 17 and Table 18, we present the quantitative comparison between DSEG-LIME ($d=2$) using EfficientNetB4 and SLIC, as reported in the main paper. The hyperparameter settings were consistent across the evaluations, except for compactness. We established a minimum threshold of 0.05 for values to mitigate the impact of poor segmentation performance, which often resulted in too small segments. Additional segments were utilized to meet this criterion for scenarios with suboptimal segmentation. However, this compactness constraint was not applied to DSEG with depth two since its hierarchical approach naturally yields smaller and more detailed explanations, evident in Table 18. The hierarchical segmentation of $d=2$ slightly impacts stability, yet the method continues to generate meaningful explanations, as indicated by other metrics. Although our method demonstrated robust performance, it required additional time because the feature attribution process was conducted twice.
@@
Table 18: Quantitative summary - numbers depth two. The table showcases the numeric values in the same manner as in Table 17 but for numeric values.
@@
Exemplary explanations. DSEG-LIME introduces a hierarchical feature generation approach, allowing users to specify segmentation granularity via tree depth. Figure 6 displays five examples from our evaluation, with the top images showing DSEG’s explanations at a hierarchy depth of one and the bottom row at a depth of two. These explanations demonstrate that deeper hierarchies focus on smaller regions. However, the banana example illustrates a scenario where no further segmentation occurs if the concept, like a banana, lacks sub-components for feature generation, resulting in identical explanations at both depths.
Figure 6: DSEG depth two. The figure displays exemplary images from the evaluation dataset, illustrating DSEG explanations at $d=2$ of hierarchical segmentation. These images serve as complementary examples to the paper’s discussion on the projectile, enhancing the illustration of the concept.
In Figure 7, another instance is explained with DSEG and $d=2$, showing a black-and-white image of a projectile. Here, we see the corresponding explanation for each stage, starting with the first iteration with the corresponding segmentation map. In the second iteration, we see the segment representing the projectile split into its finer segments - the children nodes of the parent node - with the corresponding explanation below.
@@
Case study. We examine the case presented in Figure 3, where DSEG initially segments the image into various layers with overlapping features, establishing a segmentation hierarchy through composition. In the first iteration, LIME focuses solely on the segments just beneath the root node - the parent segments that cannot be merged into broader concepts. From this segmentation map, LIME determines the feature importance scores, identifying the airplane as the most crucial element in the image. In the subsequent iteration, illustrated in Figure 8, DSEG generates an additional segmentation map that further divides the airplane into finer components for detailed analysis. The explanation in this phase emphasizes the airplane’s body, suggesting that this concept of the ’Airliner’ is most significant.
Figure 8: Airliner explanation with depth two. The same example as in Figure 3 but with segmentation hierarchy of two for the explanation. This example includes the children nodes of the most significant parent node in the segmentation map for feature importance calculation.
Appendix C Further experiments with DSEG-LIME
@@ C.1 Ablation study
For the ablation study, we examine how the number of segments evolves across different stages of the segmentation process as we vary the threshold for removing segments smaller than the hyperparameter $\theta$ (with values [100, 300, 500, 1000, 2000]). Additionally, we assess the behavior of empty spaces within the segmented regions across all images in the dataset. The analysis focuses on three key points: the number of segments immediately after the initial automated segmentation, after hierarchical sorting, and after the removal of undersized segments, following the complete DSEG approach. The empty space is evaluated before it is filled with adjacent segments. A comprehensive overview of the metrics for these steps is presented in Figure 9.
@@
C.2 Explaining wrong classification
Here, we explore how DSEG can aid in explaining a model’s misclassification. Unlike the previous analysis in Section 5.2.1, where metrics were assessed under simulations involving a model with randomized weights (Random Model) or random predictions (Random Expl.), this case focuses on a real misclassification by EfficientNetB4, free from external manipulation. This allows for a more genuine examination of DSEG’s ability to explain incorrect classifications under normal operating conditions.

### v5

@@ 5.3 Qualitative evaluation
User study. Following the methodology by Chromik and Schuessler [12], we conducted a user study (approved by the institute’s ethics council) to assess the interpretability of the explanations. This study involved 87 participants recruited via Amazon Mechanical Turk (MTurk) and included 20 random images in our dataset (Section˜D.1). These images were accompanied by explanations using DSEG and other segmentation techniques within the LIME framework (Section˜5.1).
@@
Participants rated the explanations on a scale from 1 (least effective) to 5 (most effective) based on their intuitive understanding and the predicted class. Table˜3 summarizes the average scores, the cumulative number of top-rated explanations per instance, and the statistical significance of user study results for each segmentation approach. DSEG is most frequently rated as the best and consistently ranks high, even when it is not the leading explanation. Paired t-tests indicate that DSEG is statistically significantly superior (additional results in Section˜D.2).
Exemplary Explanations. Figure˜4 displays five examples from our evaluation, with the top images showing DSEG’s explanations at a hierarchy depth of one and the bottom row at a depth of two. These explanations demonstrate that deeper hierarchies focus on smaller regions. However, the banana example illustrates a scenario where no further segmentation occurs if the concept, like a banana, lacks sub-components for feature generation, resulting in identical explanations at both depths. An example with a depth of three applied to a medical CT image is presented in Section˜C.2.
Figure 4: DSEG examples. Exemplary images from the evaluation dataset, illustrating DSEG explanations at $d=1$, and $d=2$.
These five images (with $d=1$) were also included in the user study. For the complete set of 100 evaluation images and their attribution maps, please refer to the supplementary material (Section˜C.7).
@@ 5.4 Limitations and future work
DSEG-LIME performs feature generation directly on images before inputting them into the model for explanation. For models like ResNet with smaller input sizes [22], the quantitative advantages are less evident (Section˜B.1). Experiments have shown that better results can be achieved with a lower SAM stability score threshold. Furthermore, substituting superpixels with a specific value in preservation and deletion evaluations can introduce an inductive bias [20]. To reduce this bias, using a generative model to synthesize replacement areas could offer a more neutral alteration. Additionally, future work should thoroughly evaluate feature attribution maps to ensure that methods assign significant attributions to the correct regions, as in Section˜C.7. This comprehensive assessment is essential to verify the interpretability and reliability of such methods. Lastly, our approach, like any other LIME-based method [49, 54, 60, 61], does not assume a perfect match between the explanation domains and the model’s actual domains since it simplifies the model by a local surrogate. Nonetheless, our quantitative analysis confirms that the approximations closely reflect the model’s behavior. Future work could focus on integrating foundation models directly into the system through a model-intrinsic approach, similar to [52].
No free lunch. While DSEG shows promising results across various domains, it is not universally applicable—its performance depends on the underlying segmentation model. If the model fails, for instance, due to a lack of domain-specific knowledge or the complexity of the feature generation task, the resulting concept extraction can be suboptimal (Section˜C.6). To address this limitation, future work could explore alternative segmentation backbones, such as HSSN [56] or HIPPIE [30], in place of SAM (or DETR). Moreover, recent extensions of SAM have introduced domain-adaptive modules that can improve segmentation in specialized contexts [11].
@@ 6 Conclusion
In this study, we introduced DSEG-LIME, an extension to the LIME framework, incorporating segmentation foundation models for feature generation with hierarchical feature calculation. This approach ensures that the generated features more accurately reflect human-recognizable concepts, enhancing the interpretability of explanations. Furthermore, we refined the process of feature attribution within LIME through an iterative method, establishing a segmentation hierarchy that contains the relationships between components and their subcomponents. In Section˜C.3, we show that our idea also helps explain a model’s wrong classifications. Through a comprehensive two-part evaluation, split into quantitative and qualitative analyses, DSEG emerged as the superior method, outperforming other LIME-based approaches in most evaluated metrics. The adoption of foundational models marks a significant step toward enhancing post-hoc and model-agnostic interpretability of deep learning models.
{ack}
This research was supported in part by the German Federal Ministry for Economic Affairs and Climate Action of Germany (BMWK), and in part by the German Federal Ministry for Research, Technology, and Space (BMFTR).
@@ A.1 DSEG-LIME algorithm
1:Input: $f$ (black-box model), $\zeta$ (segmentation function), $x$ (input instance), $g$ (interpretable model), $d$ (maximum depth), hphp (segmentation hyperparameters), $\theta$ (minimum segment size), $k$ (top segments to select)
@@
[…]
@@
[…]
@@
First, we identify the top-level segments, which do not occur as subparts of any other segments. These segments serve as the highest-level nodes in the hierarchical graph. Starting from these top-level segments, we apply a top-down approach to identify child segments recursively. For each parent segment, we check for segments within it; these segments are designated as child nodes of the parent in the graph. Only the selected nodes are exclusively used to train the surrogate model.
@@ A.2 Hierarchical composition
The algorithm, designed for hierarchical composition, constructs a graph to represent the segments of a given instance, as shown in Algorithm˜2. A key feature of this algorithm is the additional hyperparameter $t$, which determines when one segment is considered a subpart of another based on their overlap. Initially, the graph includes loops and redundant edges. To ensure it accurately represents the hierarchical relationships, the algorithm applies a series of predefined rules to tune the graph, removing unnecessary connections and simplifying its structure for the final representation.
@@ B.1 ResNet
In Table˜4, we detail the quantitative results for the ResNet-101 model, comparing our evaluation with the criteria used for EfficientNetB4 under consistent hyperparameter settings. The review encompasses a comparative analysis with EfficientNetB3, emphasizing performance under contrastive conditions. The findings substantiate the results obtained from EfficientNet, demonstrating that the LIME techniques exhibit unpredictable behavior in the presence of model noise or prediction shuffling, despite varied segmentation strategies. This suggests an inherent randomness in the model explanations. Furthermore, the results indicate that all XAI methodologies displayed performance levels that were inferior to those of EfficientNetB4. However, it is significant to note that DSEG surpassed all other methods in terms of performance.
@@
Table˜5 presents further findings of ResNet. SLIME with DSEG yields the lowest AUC for incremental deletion, whereas Quickshift and Felzenszwalb show the highest. WS produces the smallest superpixels for compactness, contrasting with DSEG’s larger ones. The stability analysis shows that all segmentations are almost at the same level, with QS being the best and GLIME the best-performing overall. Echoing EfficientNet’s review, segmentation defines runtime, with DSEG being the most time-consuming.
@@ B.2 VisionTransformer
Table˜6 provides the quantitative results for the VisionTransformer (ViT-384) model, employing settings identical to those used for EfficientNet and ResNet, with ViT processing input sizes of (384x384). The class-specific results within this table align closely with the performances recorded for the other models, further underscoring the effectiveness of DSEG. This consistency in DSEG performance is also evident in the data presented in Table˜7.
@@ B.3 ConvNext
Table˜8 and Table˜9 present the quantitative results pertaining to the ConvNext model [33], which is the last model that has been thoroughly evaluated in this study. The settings utilized in this evaluation remain consistent with those employed previously.
@@ B.4 DSEG compared to EAC
We conducted additional experiments with Explain Any Concept (EAC) [52], performing the same quantitative experiments as for DSEG, but with a reduced dataset comprising 50 images to optimize computational time (Section˜D.1). We began our evaluation by noting that EAC, unlike DSEG-LIME, cannot be applied to arbitrary models, which is a significant drawback of their method and prevents comprehensive comparisons. Thus, we compared our approach against LIME and EAC, in explaining ResNet. The results are listed in Table˜10 and Table˜11.
@@ B.5 DSEG within SLICE
Table˜12 reports the results for the DSEG framework applied to the SLICE [7] approach and conducts a comparative evaluation against Quickshift (QS) in LIME. The experiments were carried out on the minimized dataset consisting of 50 images in total. Each image was processed for 200 steps and repeated 8 times. The primary evaluation metric used was the average rank similarity (p = 0.3), which focused on positive and negative segments and ensured comparability by employing the minimum segment count of both approaches. Furthermore, a sign flip analysis was performed, measuring the number of segments that changed signs during the computation. This analysis considered all segments produced by each technique.
@@ B.6 DETR within DSEG
In Table˜13 and Table˜14, we conducted the DETR experiments within LIME. Based on prior results, we assess its performance by contrasting it with SLIC within the LIME framework, also utilizing the same condensed dataset comprising 50 images. Both experiments were configured with identical parameters, and DETR was implemented for basic panoptic segmentation.
Table 13: Quantitative summary - classes DETR. The table showcases metrics for EfficientNetB4, and DETR within DSEG. Results reported pertain solely to integrating DSEG and SLIC within the scope of the LIME frameworks examined.
@@
Table 14: Quantitative summary - numbers DETR. The table showcases the numeric values in the same manner as in Table˜13 but for numeric values.
@@
DETR demonstrates superior performance on the dataset compared to the LIME variants utilizing SLIC. Despite its efficacy, the segmentation quality of DETR was generally inferior to that of SAM, as evidenced by less compact explanations. This observation is further supported by the examples in Figure˜5. The visualizations reveal that DETR often segments images in ways that do not align with typical human-recognizable concepts, highlighting a potential limitation in its practical utility for generating explanatory segments. Moreover, DETR does not support the construction of a segmentation hierarchy, lacking the ability to produce finer and coarser segments, which diminishes its flexibility compared to methods such as SAM.
@@
B.7 SAM 2 within DSEG
In the main paper, we conducted experiments using SAM 1. In this part, we integrate SAM 2 [48] with the ’hiera_l’ backbone into the DSEG framework, applying a 0.8 stability score threshold. We also report the results for EfficientNetB4 on the dataset comprising 50 images in Table˜15 and Table˜16.
@@ B.7 SAM 2 within DSEG
Exemplary explanations. Figure˜6 presents explanations generated by DSEG using both SAM 1 and SAM 2, highlighting cases where the newer version of SAM enables DSEG to produce more meaningful and interpretable explanations. Each image includes explanations for the predicted class from EfficientNetB4. While SAM 2 shows improved segmentation in these examples, similar results can be obtained with SAM 1 by appropriately adjusting the hyperparameters for automatic mask generation.
@@ B.8 EfficientNetB4 with depth of two
In Table˜17 and Table˜18, we present the quantitative comparison between DSEG-LIME ($d=2$) using EfficientNetB4 and SLIC, as reported in the main paper. The hyperparameter settings were consistent across the evaluations, except for compactness. We established a minimum threshold of 0.05 for values to mitigate the impact of poor segmentation performance, which often resulted in too small segments. Additional segments were utilized to meet this criterion for scenarios with suboptimal segmentation. However, this compactness constraint was not applied to DSEG with depth two since its hierarchical approach naturally yields smaller and more detailed explanations, evident in Table˜18. The hierarchical segmentation of $d=2$ slightly impacts stability, yet the method continues to generate meaningful explanations, as indicated by other metrics. Although our method demonstrated robust performance, it required additional time because the feature attribution process was conducted twice.
@@
Table 18: Quantitative summary - numbers depth two. The table showcases the numeric values in the same manner as in Table˜17 but for numeric values.
@@
Exemplary explanations. In Figure˜7, an instance is explained with DSEG and $d=2$, showing a black-and-white image of a projectile. Here, we see the corresponding explanation for each stage, starting with the first iteration, with the corresponding segmentation map. In the second iteration, we see the segment representing the projectile split into its finer segments - the children nodes of the parent node - with the corresponding explanation below.
@@
Case study. We examine the case presented in Figure˜3, where DSEG initially segments the image into various layers with overlapping features, establishing a segmentation hierarchy through composition. In the first iteration, LIME focuses solely on the segments just beneath the root node - the parent segments that cannot be merged into broader concepts. From this segmentation map, LIME determines the feature importance scores, identifying the airplane as the most crucial element in the image. In the subsequent iteration, illustrated in Figure˜8, DSEG generates an additional segmentation map that further divides the airplane into finer components for detailed analysis. The explanation in this phase emphasizes the airplane’s body, suggesting that this concept of the ’Airliner’ is most significant.
Figure 8: Airliner explanation with depth two. The same example as in Figure˜3 but with segmentation hierarchy of two for the explanation. This example includes the children nodes of the most significant parent node in the segmentation map for feature importance calculation.
Appendix C Further experiments with DSEG
@@ C.1 Ablation study
Automatic mask generation. Ablating DSEG using only the automatic masks produced by foundation models presents notable challenges, particularly when applying methods like LIME. These automatically generated masks often result in many segments, some of which may overlap or leave parts of the image unsegmented. Such inconsistencies make a direct integration with LIME impractical. Consequently, our ablation in this setting is limited to comparing different foundation models for mask generation. We conducted experiments with SAM, SAM2, and DETR to evaluate their respective impact on DSEG’s performance.
Small cluster removal. For the ablation study, we examine how the number of segments evolves across different stages of the segmentation process as we vary the threshold for removing segments smaller than the hyperparameter $\theta$ (with values [100, 300, 500, 1000, 2000]). Additionally, we assess the behavior of empty spaces within the segmented regions across all images in the dataset. The analysis focuses on three key points: the number of segments immediately after the initial automated segmentation, after hierarchical sorting, and after the removal of undersized segments, following the complete DSEG approach. The empty space is evaluated before it is filled with adjacent segments. A comprehensive overview of the metrics for these steps is presented in Figure˜9.
@@
Hierarchical ordering. The effect of hierarchical ordering is already addressed through both qualitative and quantitative analyses across different depths of the explanation process. This hierarchical structure ensures no overlapping segments occur—a necessary condition for reliably computing feature attributions. As a result, an additional technical ablation isolating hierarchical ordering is not required.
Empty space removal. The presence of unsegmented regions within an image is a well-known issue in the segmentation domain. One potential solution for adapting LIME to this scenario is to assign these unsegmented areas to a dedicated segment—e.g., a “segment zero.” However, this approach introduces the challenge that such unsegmented areas are often non-contiguous and scattered across the image. From a user perspective, this fragmented behavior is undesirable and could negatively affect interpretability. To address this, we opted for a simple yet practical solution: reassigning unsegmented pixels to the nearest segmented region using a nearest-neighbor approach. While effective for our purposes, this strategy opens avenues for future work, where more sophisticated treatments of unsegmented areas could be explored.
C.2 Medical scenario with depth of three
Figure 10: Medical use case. Demonstration of DSEG on a medical use case using CT scan imagery. From left to right: the original CT image, the raw segmentation mask, and hierarchical segmentations generated by DSEG at depths 1, 2, and 3. Each subsequent depth introduces finer segmentation details important for explanation. This showcases the general motivation for hierarchical explanations and the applicability of DSEG in medical imaging contexts.
We demonstrate the application of our framework on a medical use case, processing CT imagery of the human body, as shown in Figure˜10. In this example, we present the original image alongside the raw segmentation mask and DSEG’s hierarchical compositions at depths 1 to 3, where each level introduces progressively finer segments relevant for explanation.
For this demonstration, we utilize the basic SAM variant. However, improved segmentation quality could be achieved by using fine-tuned models specifically adapted to the medical domain. It is important to note that we do not provide a detailed explanation of a specific instance here; rather, we aim to illustrate the general motivation behind DSEG. Thus, the presented hierarchy serves demonstration purposes but highlights the feasibility and applicability of our framework when paired with appropriate segmentation models for medical data.
Furthermore, we show that the segmentation hierarchy can extend beyond depth 2, as previously evaluated in Section˜B.8. As a final remark, if a user requests an explanation deeper than the maximum available hierarchy (determined by the most activated nodes), the explanation terminates due to the inherent limitation in segmentation depth.
C.3 Explaining wrong classification
Here, we explore how DSEG can aid in explaining a model’s misclassification. Unlike the previous analysis in Section˜5.2.1, where metrics were assessed under simulations involving a model with randomized weights (Random Model) or random predictions (Random Expl.), this case focuses on a real misclassification by EfficientNetB4, free from external manipulation. This allows for a more genuine examination of DSEG’s ability to explain incorrect classifications under normal operating conditions.